OpenStudy (anonymous):

Solve the absolute value inequality. Express the answer using interval notation. 9 − |2x + 7| ≤ 6

OpenStudy (jdoe0001):

\(\bf 9 - |2x + 7| \le 6 \implies -9+9 - |2x + 7| \le 6-9\implies |2x + 7| \le -3\\ \quad \\ |2x + 7| \le -3\implies \begin{cases} +(2x + 7) \le -3\\ \quad \\ \bf -(2x + 7) \le -3 \end{cases}\) that'd give you 2 values for "x", one for each scenario

OpenStudy (anonymous):

But an absolute value cannot be less than a negative number, since it is constantly positive.

OpenStudy (jdoe0001):

hmmm I made a quick mistake... kinda ate the minus so \(\bf 9 - |2x + 7| \le 6 \implies -9+9 - |2x + 7| \le 6-9\implies -|2x + 7| \le -3\\ \quad \\ \textit{recall that whenever multiplying by a negative value you have to }\\ {\color{red}{ flip}}\textit{ the inequality sign}\\ \quad \\ |2x+7|\ge 3 \implies |2x + 7| \ge 3\implies \begin{cases} +(2x + 7) \ge -3\\ \quad \\ \bf -(2x + 7) \ge -3 \end{cases}\)

OpenStudy (jdoe0001):

hmm shoot.. .anyhow.. even more typos... one sec \(\bf 9 - |2x + 7| \le 6 \implies -9+9 - |2x + 7| \le 6-9\implies -|2x + 7| \le -3\\ \quad \\ \textit{recall that whenever multiplying by a negative value you have to }\\ {\color{red}{ flip}}\textit{ the inequality sign}\\ \quad \\ |2x+7|\ge 3 \implies |2x + 7| \ge 3\implies \begin{cases} +(2x + 7) \ge 3\\ \quad \\ \bf -(2x + 7) \ge 3 \end{cases}\) so.... let's try to do the 1st one

OpenStudy (jdoe0001):

\(\bf +(2x + 7) \ge -3\implies 2x+7\ge 3\) so... what should we do to isolate the "x"?

OpenStudy (jdoe0001):

ack... typos are killing me anyhow \(\bf +(2x + 7) \ge 3\implies 2x+7\ge 3\)

OpenStudy (jdoe0001):

you'd do it more or less in the same way you'd solve a linear equation, as if you were using

OpenStudy (jdoe0001):

as if you were using \(\bf 2x+7= 3\)